A garden has an iside diameter of 1.6cm. the nozzle opening has a diameter of 0.43cm/s. what is the flow speed through the nozzle?
1
Expert's answer
2020-12-21T10:35:53-0500
"\frac{dV}{dt}=Av=0.25\pi d^2v\\\frac{dV}{dt}=0.25\pi (1.6)^2(0.43)=0.865\frac{cm^3}{s}"